# Size Winding Mechanism beyond Maximum Chaos

### v1

@@ Abstract
The concept of information scrambling elucidates the dispersion of local information in quantum many-body systems, offering insights into various physical phenomena such as wormhole teleportation. This phenomenon has spurred extensive theoretical and experimental investigations. Among these, the size-winding mechanism emerges as a valuable diagnostic tool for optimizing signal detection. In this Letter, we establish a computational framework for determining the winding size distribution in large-$N$ quantum systems with all-to-all interactions, utilizing the scramblon effective theory. We obtain the winding size distribution for the large-$q$ SYK model across the entire time domain. Notably, we unveil that the manifestation of size winding results from a universal phase factor in the scramblon propagator, highlighting the significance of the Lyapunov exponent. These findings contribute to a sharp and precise connection between operator dynamics and the phenomenon of wormhole teleportation.
@@
In other words, the phase of the operator wavefunction depends solely on the size $n$ through a linear function. We provide a concise review of the relationship between the teleportation signal and size winding in the supplementary material [18] for completeness. The size winding can also be probed by combining the standard size distribution $P(n,t)$ with the winding size distribution $Q(n,t)$, defined as:
$P(n,t)$ $=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left|c_{j_{1}j_{2}\cdots j_{n}}(t)\right|^{2},$ (2)
$Q(n,t)$ $=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left(c_{j_{1}j_{2}\cdots j_{n}}(t)\right)^{2}.$
The perfect size winding in Eq. (1) is then equivalent to having $Q(n,t)/P(n,t)=e^{i(a+bn)}$. Known examples of systems with (near-)perfect size winding include the large-$q$ SYK model in the early-time regime and holographic systems with semi-classical gravity [15, 16, 14]. In this letter, we present a refined understanding of information scrambling by computing the size distribution $P(n,t)$ and the winding size distribution $Q(n,t)$ in large-$N$ quantum mechanics using the scramblon effective theory [19, 6, 20, 21]. We show that the scramblon propagator contributing to the generating function of $Q(n,t)$ acquires a pure imaginary time shift compared to $P(n,t)$, which is the origin of the size winding. This does not necessarily depend on maximal chaos, but rather relies on the universal chaotic behavior of the system [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 2, 3, 36, 37]. Applying this approach to the large-$q$ SYK model [38, 39, 40, 41] yields the size winding distribution for the full range of time, thereby extending existing results [15, 16, 14]. The results (illustrated in FIG. 1) reveal two different regimes for size winding. The first regime is the early-time regime $\varkappa^{-1}\ll t\ll t_{\text{sc}}$ with scrambling time $t_{\text{sc}}$ in the large-$N$ limit, where typical operators have $n\sim\mathcal{O}(1)$. The second regime corresponds to the long-time limit $t-t_{\text{sc}}\gg\varkappa^{-1}$, where finite-$N$ corrections become crucial, indicating a completely different mechanism for size winding.
@@
Generating functions.– To cover the full time range in which typical operators can get scrambled across the entire system, we normalize $s=n/N$ and consider the thermodynamic limit $N\rightarrow\infty$ [6]. Consequently, $s\in[0,1]$ becomes a continuous variable and we further define continuous distributions $\mathcal{P}(s,t)=NP(sN,t)$ and $\mathcal{Q}(s,t)=NQ(sN,t)$. When computing distribution functions, it is more convenient to introduce generating functions through a Laplace transform. We have
@@
and similarly for $\mathcal{S}_{\mathcal{P}}$. All moments of $s$ can be computed by taking derivatives with respect to $\nu$. After obtaining the generating functions, the size distribution and winding size distribution can be determined through an inverse Laplace transform. Ref. [5] shows that the generating function of the size distribution can be naturally described by a correlation function in a double system. Subsequently, this connection has been extended to spin models [21] and the winding size distribution [15, 16, 14]. We first introduce an auxiliary system with $N$ Majorana fermions $\psi_{j}$. We prepare the double system in the EPR state specified by $(\chi_{j}+i\psi_{j})|\text{EPR}\rangle=0$, which is a vacuum state for complex fermion modes $c_{j}=(\chi_{j}+i\psi_{j})/2$. Under this choice, applying a string operator of Majorana fermions $\chi_{j_{1}}\chi_{j_{2}}\ldots\chi_{j_{n}}$ to the EPR state creates $n$ complex fermions. Therefore, computing the size of operators is mapped to counting the particle number $\hat{n}=\sum_{j}c^{\dagger}_{j}c_{j}=\sum_{j}{(1+i\chi_{j}\psi_{j})}/{2}$. This leads to
$\mathcal{S}_{\mathcal{P}}=\langle\text{EPR}|\rho^{1/2}\chi_{k}(t)e^{-\frac{\nu}{2N}\sum_{j}(1+i\chi_{j}\psi_{j})}\chi_{k}(t)\rho^{1/2}|\text{EPR}\rangle,$ (4)
$\mathcal{S}_{\mathcal{Q}}=\langle\text{EPR}|\rho^{1/2}\chi_{k}(t)e^{-\frac{\nu}{2N}\sum_{j}(1+i\chi_{j}\psi_{j})}\rho^{1/2}\chi_{k}(t)|\text{EPR}\rangle,$
@@
where we omit the arguments for conciseness. The generalization to spin models is straightforward using the size-total spin correspondence introduced in [21]. Both generating functions possess a path-integral representation on the double Keldysh contour [42]. A closer examination reveals an important distinction in their imaginary time configurations, as illustrated in FIG 2. Keeping to the first order in $1/N$ [18] (reasons for going beyond the zeroth order are explained later), both $\mathcal{S}_{\mathcal{P}}(\nu,t)$ and $\mathcal{S}_{\mathcal{Q}}(\nu,t)$ can be expressed in a unified form
$\mathcal{S}_{\mathcal{P}/\mathcal{Q}}=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\left<\text{T}_{c}\chi_{k}(T_{1})\chi_{k}(T_{2})\prod_{j}{\Big{(}1+\frac{\nu~{}\chi_{j}(T_{3})\chi_{j}(T_{4})}{2N}\Big{)}}\right>.$ (5)
@@
Here, $\text{T}_{c}$ represents the contour ordering operator. We introduce $T=t-{i\mkern 1.0mu}\tau$, where $t$ represents real-time and $\tau$ represents imaginary time. Both the traditional size distribution and the winding size distribution contain out-of-time-order (OTO) correlations [22, 2, 3, 36, 37]. Therefore, the typical timescale for the evolution of the distribution functions is the scrambling time, and their calculation should involve recent developments in the scramblon effective theory [19, 6, 20, 21]. Moreover, in comparison to $\mathcal{S}_{\mathcal{P}}(\nu,t)$, there is a $\beta/2$ imaginary time shift for $T_{2}$ in the winding size generation function, which will give rise to the size winding phenomena. Scramblon calculation.– In the short-time limit specified by $\varkappa t\lesssim 1$, the evolution involves contributions from all microscopic details, making it non-universal. Therefore, our primary interest lies in the universal physics for $\varkappa t\gg 1$, where OTO-correlations dominate. In the scramblon effective theory, operators respect the Wick theorem unless they manifest OTO correlations, which are mediated by collective modes known as scramblons [41, 43]. To warm up, a four-point OTO correlator (OTOC) can be computed by summing up diagrams where two pairs of operators interact by exchanging an arbitrary number of scramblons denoted by $m$:
@@
where $\lambda=C^{-1}e^{{i\mkern 1.0mu}\varkappa\frac{\beta}{4}+\varkappa\frac{T_{1}+T_{2}-T_{3}-T_{4}}{2}}$ is the propagator of the scrambling modes. The crucial phase factor $e^{{i\mkern 1.0mu}\varkappa{\beta}/{4}}$ ensures that the result is real for an equally spaced case with $\tau_{1}-\tau_{3}=\tau_{2}-\tau_{4}=\beta/4$. For the size generating function, configurations in FIG 2 corresponds to a real $\lambda=C^{-1}e^{\varkappa t}\equiv\lambda_{0}$ while for the winding size distribution, we find $\lambda=e^{i\varkappa\beta/4}\lambda_{0}$ due to the additional imaginary time shift. The prefactor $C$ is of the order of $N$ and, consequently, suppresses higher-order terms at an early time. We also introduce the scattering amplitude between fermions and $m$ scramblons in the future or past as $\Upsilon^{{\mathrm{R}},m}$ or $\Upsilon^{{\mathrm{A}},m}$. Applying a similar treatment, the generating functions $\mathcal{S}_{\mathcal{P}}$ and $\mathcal{S}_{\mathcal{Q}}$ in Eq. (5) can be computed by summing over all configurations of scramblons. Furthermore, following the manipulations in Ref. [19], the vertex functions $\Upsilon^{{\mathrm{R}},m}$ and $\Upsilon^{{\mathrm{A}},m}$ can be expressed as moments of $h^{{\mathrm{R}}/{\mathrm{A}}}$, i.e., $\Upsilon^{{\mathrm{R}}/{\mathrm{A}},m}(T)=\int_{0}^{+\infty}y^{m}h^{{\mathrm{R}}/{\mathrm{A}}}(y,T)\mathop{}\!\mathrm{d}y$, which describe the distribution of the perturbation strength generated by Majorana fermion operators. Leaving details into the supplementary material [18], we obtain closed-form results:
@@
$\begin{split}\mathcal{S}_{\mathcal{P}/\mathcal{Q}}&=\int\mathop{}\!\mathrm{d}y\,h^{\mathrm{R}}(y,T_{12})e^{\frac{\nu^{2}}{8N}(1-f^{A}(\lambda y,-i\beta/2)^{2})-\frac{\nu}{2}(1-f^{A}(\lambda y,-i\beta/2))},\\ \end{split}$ (7)
@@
The function $f^{A}$ represents the Laplace transform of $h^{{\mathrm{A}}}$, $f^{\mathrm{A}}(x,T)=\int_{0}^{+\infty}e^{-xy}h^{{\mathrm{A}}}(y,T)~{}\mathop{}\!\mathrm{d}y$. It describes the fermion two-point function under the influence of scramblon perturbation with strength $x$. By applying the inverse Laplace transform, we can approximate the size distribution or winding size distribution as
@@
with $\sigma^{2}={(1-f^{A}(\lambda y,-i\beta/2)^{2})}/{4N}$. Eq. (8) is the central result of this work, which is valid for generic chaotic large-$N$ quantum systems with all-to-all interactions, beyond models with maximal chaos. In the following sections, we analyze this result to unveil regimes which exhibits size winding. Size winding at large-$N$.– We can divide the the evolution into three time regimes: the early-time regime with $1\ll\varkappa t\ll t_{\text{sc}}$, the intermediate time regime with $t\approx t_{\text{sc}}$, and the long-time limit with $t-t_{\text{sc}}\gg\varkappa^{-1}$. In the first two regimes, the operator size distribution has an $\mathcal{O}(1)$ variance in continuous size $s$, allowing us to replace the Gaussian function in Eq. (8) with a delta function, as in the standard large-$N$ limit [6]. In this case, the result can be expressed as

### v2

# Size Winding Mechanism beyond Maximal Chaos

@@ Abstract
The concept of information scrambling elucidates the dispersion of local information in quantum many-body systems, offering insights into various physical phenomena such as wormhole teleportation. This phenomenon has spurred extensive theoretical and experimental investigations. Among these, the size-winding mechanism emerges as a valuable diagnostic tool for optimizing signal detection. In this work, we establish a computational framework for determining the winding size distribution in large-$N$ quantum systems with all-to-all interactions, utilizing the scramblon effective theory. We obtain the winding size distribution for the large-$q$ SYK model across the entire time domain. Notably, we unveil that the manifestation of size winding results from a universal phase factor in the scramblon propagator, highlighting the significance of the Lyapunov exponent. These findings contribute to a sharp and precise connection between operator dynamics and the phenomenon of wormhole teleportation.
1 Introduction
During chaotic unitary evolution, localized initial information within interacting many-body systems rapidly disseminates across the entire system—a phenomenon known as information scrambling [1, 2]. In the Heisenberg picture, this scrambling process is elucidated through the growth of simple operators, typically quantified by the operator size distribution [3, 4, 5, 6], which is a probability distribution in a specific operator basis. Nevertheless, a comprehensive understanding of operator growth entails considering both the amplitude and the phase of the generic operator wavefunction. In particular, the phase is crucial for several fascinating and counterintuitive properties of quantum systems, including wormhole teleportation [7, 8, 9, 10, 11, 12, 11, 13, 14, 15, 16], which has experimental investigation on the Google Sycamore processor with nine qubits [17], indicating an exciting era of studying “quantum gravity in lab”. Previous investigations have identified the size-winding mechanism as a promising candidate [15, 16, 17]. In this proposal, to maximize the teleportation signal, the phase of the operator wavefunction should satisfy a specific pattern, as we will now elaborate.
@@ 1 Introduction
In other words, the phase of the operator wavefunction depends solely on the size $n$ through a linear function. We provide a concise review of the relationship between the teleportation signal and size winding in the Appendix A for completeness. The size winding can also be probed by combining the standard size distribution $P(n,t)$ with the winding size distribution $Q(n,t)$, defined as:
$P(n,t)=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left|c_{j_{1}j_{2}\cdots j_{n}}(t)\right|^{2},\qquad Q(n,t)=\sum_{j_{1}<j_{2}<\cdots<j_{n}}\left(c_{j_{1}j_{2}\cdots j_{n}}(t)\right)^{2}.$ (2)
The perfect size winding in Eq. (1) is then equivalent to having $Q(n,t)/P(n,t)=e^{i(a+bn)}$. Known examples of systems with (near-)perfect size winding include the large-$q$ SYK model in the early-time regime and holographic systems with semi-classical gravity [15, 16, 14].
In this work, we present a refined understanding of information scrambling by computing the size distribution $P(n,t)$ and the winding size distribution $Q(n,t)$ in large-$N$ quantum mechanics using the scramblon effective theory [18, 6, 19, 20]. We show that the scramblon propagator contributing to the generating function of $Q(n,t)$ acquires a pure imaginary time shift compared to $P(n,t)$, which is the origin of the size winding. This does not necessarily depend on maximal chaos, but rather relies on the universal chaotic behavior of the system [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 2, 3, 35, 36]. Applying this approach to the large-$q$ SYK model [37, 38, 39, 40] yields the size winding distribution for the full range of time, thereby extending existing results [15, 16, 14]. The results (illustrated in FIG. 1) reveal two different regimes for size winding. The first regime is the early-time regime $\varkappa^{-1}\ll t\ll t_{\text{sc}}$ with scrambling time $t_{\text{sc}}$ in the large-$N$ limit, where typical operators have $n\sim\mathcal{O}(1)$. The second regime corresponds to the long-time limit $t-t_{\text{sc}}\gg\varkappa^{-1}$, where finite-$N$ corrections become crucial, indicating a completely different mechanism for size winding.
[…]
@@
2 General Analysis
To cover the full time range in which typical operators can get scrambled across the entire system, we normalize $s=n/N$ and consider the thermodynamic limit $N\rightarrow\infty$ [6]. Consequently, $s\in[0,1]$ becomes a continuous variable and we further define continuous distributions $\mathcal{P}(s,t)=NP(sN,t)$ and $\mathcal{Q}(s,t)=NQ(sN,t)$. When computing distribution functions, it is more convenient to introduce generating functions through a Laplace transform. We have
@@ 2 General Analysis
and similarly for $\mathcal{S}_{\mathcal{P}}$. All moments of $s$ can be computed by taking derivatives with respect to $\nu$. After obtaining the generating functions, the size distribution and winding size distribution can be determined through an inverse Laplace transform.
2.1 Generating functions
@@
Ref. [5] shows that the generating function of the size distribution can be naturally described by a correlation function in a double system. Subsequently, this connection has been extended to spin models [20] and the winding size distribution [15, 16, 14]. We first introduce an auxiliary system with $N$ Majorana fermions $\psi_{j}$. We prepare the double system in the EPR state specified by $(\chi_{j}+i\psi_{j})|\text{EPR}\rangle=0$, which is a vacuum state for complex fermion modes $c_{j}=(\chi_{j}+i\psi_{j})/2$. Under this choice, applying a string operator of Majorana fermions $\chi_{j_{1}}\chi_{j_{2}}\ldots\chi_{j_{n}}$ to the EPR state creates $n$ complex fermions. Therefore, computing the size of operators is mapped to counting the particle number $\hat{n}=\sum_{j}c^{\dagger}_{j}c_{j}=\sum_{j}{(1+i\chi_{j}\psi_{j})}/{2}$. This leads to
$\mathcal{S}_{\mathcal{P}}(\nu,t)=\langle\text{EPR}|\rho^{1/2}\chi_{k}(t)e^{-\frac{\nu}{2N}\sum_{j}(1+i\chi_{j}\psi_{j})}\chi_{k}(t)\rho^{1/2}|\text{EPR}\rangle,$ (4)
$\mathcal{S}_{\mathcal{Q}}(\nu,t)=\langle\text{EPR}|\rho^{1/2}\chi_{k}(t)e^{-\frac{\nu}{2N}\sum_{j}(1+i\chi_{j}\psi_{j})}\rho^{1/2}\chi_{k}(t)|\text{EPR}\rangle,$
where we omit the arguments for conciseness. The generalization to spin models is straightforward using the size-total spin correspondence introduced in [20].
@@
Both generating functions possess a path-integral representation on the double Keldysh contour [41]. A closer examination reveals an important distinction in their imaginary time configurations, as illustrated in FIG 2. Keeping to the first order in $1/N$ (reasons for going beyond the zeroth order are explained later), both $\mathcal{S}_{\mathcal{P}}(\nu,t)$ and $\mathcal{S}_{\mathcal{Q}}(\nu,t)$ can be expressed in a unified form
$\mathcal{S}_{\mathcal{P}/\mathcal{Q}}(\nu,t)=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\left<\text{T}_{c}\chi_{k}(T_{1})\chi_{k}(T_{2})\prod_{j}{\Big{(}1+\frac{\nu~{}\chi_{j}(T_{3})\chi_{j}(T_{4})}{2N}\Big{)}}\right>.$ (5)
@@
Here, $\text{T}_{c}$ represents the contour ordering operator. We introduce $T=t-{i\mkern 1.0mu}\tau$, where $t$ represents real-time and $\tau$ represents imaginary time. Both the traditional size distribution and the winding size distribution contain out-of-time-order (OTO) correlations [21, 2, 3, 35, 36]. Therefore, the typical timescale for the evolution of the distribution functions is the scrambling time, and their calculation should involve recent developments in the scramblon effective theory [18, 6, 19, 20]. Moreover, in comparison to $\mathcal{S}_{\mathcal{P}}(\nu,t)$, there is a $\beta/2$ imaginary time shift for $T_{2}$ in the winding size generation function, which will give rise to the size winding phenomena.
2.2 Scramblon calculation
In the short-time limit specified by $\varkappa t\lesssim 1$, the evolution involves contributions from all microscopic details, making it non-universal. Therefore, our primary interest lies in the universal physics for $\varkappa t\gg 1$, where OTO-correlations dominate. In the scramblon effective theory, operators respect the Wick theorem unless they manifest OTO correlations, which are mediated by collective modes known as scramblons [40, 42]. To warm up, a four-point OTO correlator (OTOC) can be computed by summing up diagrams where two pairs of operators interact by exchanging an arbitrary number of scramblons denoted by $m$:
@@ 2.2 Scramblon calculation
where $\lambda=C^{-1}e^{{i\mkern 1.0mu}\varkappa\frac{\beta}{4}+\varkappa\frac{T_{1}+T_{2}-T_{3}-T_{4}}{2}}$ is the propagator of the scrambling modes. The crucial phase factor $e^{{i\mkern 1.0mu}\varkappa{\beta}/{4}}$ ensures that the result is real for an equally spaced case with $\tau_{1}-\tau_{3}=\tau_{2}-\tau_{4}=\beta/4$. For the size generating function, configurations in FIG 2 corresponds to a real $\lambda=C^{-1}e^{\varkappa t}\equiv\lambda_{0}$ while for the winding size distribution, we find $\lambda=e^{i\varkappa\beta/4}\lambda_{0}$ due to the additional imaginary time shift. The prefactor $C$ is of the order of $N$ and, consequently, suppresses higher-order terms at an early time. We also introduce the scattering amplitude between fermions and $m$ scramblons in the future or past as $\Upsilon^{{\mathrm{R}},m}$ or $\Upsilon^{{\mathrm{A}},m}$.
Applying a similar treatment, the generating functions $\mathcal{S}_{\mathcal{P}}$ and $\mathcal{S}_{\mathcal{Q}}$ in Eq. (5) can be computed by summing over all configurations of scramblons. We expand the product of Majorana operators to the Majorana string basis, which leads to
$\begin{split}\prod_{j=1}^{N}{\left(1+\frac{\nu}{2N}\chi_{j}(T_{3})\chi_{j}(T_{4})\right)}&=\sum_{n=1}^{N}\sum_{|\bm{J}|=n}\left((-1)^{\frac{(n-1)n}{2}}\chi_{\bm{J}}(T_{3})\chi_{\bm{J}}(T_{4})\left(\frac{\nu}{2N}\right)^{n}\right)\\ \end{split}.$ (7)
Our approach to compute the generating function is analogous to the calculation of the OTOC as a summation of scramblon modes. In this case, we find
$\mathcal{S}_{\mathcal{P}/\mathcal{Q}}(\nu,t)=$ $e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\sum_{n=0}^{N}\sum_{|\bm{J}|=n}\sum_{m_{1},\cdots,m_{n}}\Upsilon^{{\mathrm{R}},\sum_{l=1}^{n}m_{l}}(T_{12})$ (8)
×$\times\left(\frac{(-\lambda)^{m_{1}}}{m_{1}!}\Upsilon^{{\mathrm{A}},m_{1}}(T_{34})\cdots\frac{(-\lambda)^{m_{n}}}{m_{n}!}\Upsilon^{{\mathrm{A}},m_{n}}(T_{34})\right)\left(\frac{\nu}{2N}\right)^{n}.$
Following the manipulations in Ref. [18], the vertex functions $\Upsilon^{{\mathrm{R}},m}$ and $\Upsilon^{{\mathrm{A}},m}$ can be expressed as moments of $h^{{\mathrm{R}}/{\mathrm{A}}}$, i.e., $\Upsilon^{{\mathrm{R}}/{\mathrm{A}},m}(T)=\int_{0}^{+\infty}y^{m}h^{{\mathrm{R}}/{\mathrm{A}}}(y,T)\mathop{}\!\mathrm{d}y$, which describe the distribution of the perturbation strength generated by Majorana fermion operators. To proceed, it is important to note that the retarded vertex in Eq. (8) can be expressed through the function $h^{R}$ as
$\Upsilon^{{\mathrm{R}},\sum_{l}m_{l}}(T_{12})=\int_{0}^{\infty}\mathop{}\!\mathrm{d}yy^{\sum_{l}m_{l}}h^{R}(y,T_{12}),$ (9)
and we can combine the introduced variable $y$ to scramblon propagator $\lambda$ and the advanced vertex to get another function $f^{A}$, which is defined as
$\sum_{l}\frac{(-\lambda y)^{m_{l}}}{m_{l}!}\Upsilon^{{\mathrm{A}},m_{l}}(T_{34})=\int_{0}^{\infty}\mathop{}\!\mathrm{d}y_{l}e^{-\lambda yy_{l}}h^{A}(y_{l},T_{34})=f^{A}(\lambda y,T_{34}).$ (10)
Physically, $f^{A}$ characterizes the fermion two-point function influenced by scramblon perturbations. Therefore, we arrive at
@@
$\begin{split}\mathcal{S}_{\mathcal{P}/\mathcal{Q}}(\nu,t)&=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\int_{0}^{\infty}\mathop{}\!\mathrm{d}yh^{R}(y,T_{12})\sum_{n=0}^{N}\sum_{|\bm{J}|=n}\left(f^{A}(\lambda y,T_{34})\right)^{n}\left(\frac{\nu}{2N}\right)^{n}\\ &=e^{-\frac{\nu}{2}+\frac{\nu^{2}}{8N}}\int_{0}^{\infty}\mathop{}\!\mathrm{d}yh^{R}(y,T_{12})\left(1+f^{A}(\lambda y,T_{34})\left(\frac{\nu}{2N}\right)\right)^{N}\end{split}$ (11)
Treating $\frac{\nu}{2N}$ as a small parameter, we expand the expression to the second order of $\left(\frac{\nu}{2N}\right)^{2}$ in the exponential term. We obtain closed-form results:
$\begin{split}\mathcal{S}_{\mathcal{P}/\mathcal{Q}}(\nu,t)&=\int\mathop{}\!\mathrm{d}y\,h^{\mathrm{R}}(y,T_{12})e^{\frac{\nu^{2}}{8N}(1-f^{A}(\lambda y,-i\beta/2)^{2})-\frac{\nu}{2}(1-f^{A}(\lambda y,-i\beta/2))},\\ \end{split}$ (12)
@@
It describes the fermion two-point function under the influence of scramblon perturbation with strength $x$. By applying the inverse Laplace transform, we can approximate the size distribution or winding size distribution as
@@
with $\sigma^{2}={(1-f^{A}(\lambda y,-i\beta/2)^{2})}/{4N}$. Eq. (13) is the central result of this work, which is valid for generic chaotic large-$N$ quantum systems with all-to-all interactions, beyond models with maximal chaos. In the following sections, we analyze this result to unveil regimes which exhibits size winding.
3 Size winding at large-$N$
We can divide the the evolution into three time regimes: the early-time regime with $1\ll\varkappa t\ll t_{\text{sc}}$, the intermediate time regime with $t\approx t_{\text{sc}}$, and the long-time limit with $t-t_{\text{sc}}\gg\varkappa^{-1}$. In the first two regimes, the operator size distribution has an $\mathcal{O}(1)$ variance in continuous size $s$, allowing us to replace the Gaussian function in Eq. (13) with a delta function, as in the standard large-$N$ limit [6]. In this case, the result can be expressed as